OpenStudy (anonymous):

find the angle of intersection between the curve x^2-y^2=a^2 and x^2+y^2=a^2 (sqr rtof 2)?pls help me

OpenStudy (anonymous):

\[} find the \angle of intersection \between the curve x^{2}-y^{2}=a^{2} and x^{2}+y^{2=a^{2}\]

OpenStudy (anonymous):

solution: i got slope m1= dy/dx=x/y and slope m2=dy/dx= (-x/y) and \[\Theta=\tan^{-1}(2xy/y^{2}-x^{2})\]

OpenStudy (campbell_st):

the points of intersection of the curves are at the \[\pm \sqrt{a}\]

OpenStudy (anonymous):

but how do i get ans: \[ \Theta=\tan^{-1}(\pm a^2/-a^{2})\]

OpenStudy (anonymous):

i got the dinomtr as -a^{2}

OpenStudy (campbell_st):

so are you asking about the tangents to the curves at these points... as they are vertical lines \[x = \pm a\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how do i get the numerator 2xy = pm a^{2}

OpenStudy (campbell_st):

oops just reread

OpenStudy (anonymous):

from the question we have to fine the theta

OpenStudy (anonymous):

i got theta = tan^{-1} (2xy/y^{2}-x^{2})

OpenStudy (anonymous):

\[\theta = \tan^{-1} (2xy/y^{2}-x^{2}) = \tan^{-1}(\pm a^{2}/-a^{2}\]

OpenStudy (anonymous):

i gotthe dinomntr -a^{2} from the question'\[\X^{2}-y^{2}\]\

OpenStudy (anonymous):

pls hlp me

OpenStudy (anonymous):

how do i get the numeratr \[\ pm a^{2}\]\

OpenStudy (anonymous):

\[\x^{2}+Y^{2}=a^{2}sqrt{2}\]\

OpenStudy (anonymous):

\[\(x+y)^{2}-2xy= a^{2}sqrt{2}\]\

OpenStudy (anonymous):

how to solve aftr ths to get 2xy = \[\pm a^{2}\]\

OpenStudy (anonymous):

are these your equations? x^2-y^2=a^2 and x^2+y^2=a^2

OpenStudy (anonymous):

no the question s

OpenStudy (anonymous):

i don't understand that last line in your question.

OpenStudy (anonymous):

x^2-y^2 =a^2 and x^2+y^2=a^2 sqrt{2}

OpenStudy (campbell_st):

I think you need a nominal value for a... perhaps a = 1

OpenStudy (anonymous):

first one is a hyperbola, second one is a circle with slightly bigger radius than the transverse and conjugate axis of the hyperbola. they will intersect at 4 points. and you're looking for the angle in which the tangent lines intersect at these points?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

because of symmetry, i don't think that would be too hard eh? add those two equations to get rid of the y^2.... what do you get for x?

OpenStudy (anonymous):

this s an example sum in my book it s given theta= tan^{-1} (2xy/y^2-x^2)=tan^-1(pm a^2/-a^2)

OpenStudy (anonymous):

is x =a sqrt (1+ sqrt2)/sqrt2

OpenStudy (anonymous):

sorry man, that thing you just put up don't make sense to me. i was gonna do it with finding the slope of the tangent lines at those intersections and using those slopes to calculate the angle between them.

OpenStudy (anonymous):

yes it s to find the angle btween them

OpenStudy (anonymous):

i got the slopes m1=x/y and m2=(-x/y)

OpenStudy (anonymous):

by substituting in formula theta =tan^{-1} [(m1-m2)/(1-m1m2)] i got tan^{-1}(2xy/y^2-x^2)

OpenStudy (anonymous):

sorry formula tan^{-1}[(m1-m2)/1+m1m2]

OpenStudy (anonymous):

 i got a different x-coordinate for my intercept